OpenStudy (anonymous):

What is the modulus and argument of the complex number -5i ? help is appreciated <3

OpenStudy (anonymous):

Modulus is 5 and the argument is 3pi/2

OpenStudy (jdoe0001):

\(\bf -5i\implies 0-5i\implies \begin{array}{ccllll} a&b\\ \hline\\ 0&-5i \end{array}\quad \begin{array}{llll} modulus\implies r=\sqrt{a^2+b^2}\\ \quad \\ \bf argument\implies \theta=tan^{-1}\left(\frac{b}{a}\right) \end{array}\)

OpenStudy (jdoe0001):

the "b" of course, doesn't include the 'i'

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Anytime babe!